OpenStudy (anonymous):

How many ways can you purchase 5 CDs if there are 6 to choose from, 2 cassettes if there are 5 to choose from, and 4 DVDs if there are 8 to choose from?

OpenStudy (elonasushchik):

(65)×(52)×(84) is how you do this

OpenStudy (elonasushchik):

(65)=6 (52)=5×42=10 (84)=8×7×6×54×3×2=2×7×5=70

OpenStudy (anonymous):

You are amazing omg thank you

OpenStudy (elonasushchik):

Anytime!!!!